# A Backdoor Approach with Inverted Labels Using Dirty Label-Flipping Attacks

### v1

@@ II Related work
$\mathbf{\mathcal{B}}^{*}\in\underset{\mathbf{\mathcal{B}}\in\{0,1\}^{N},\|\mathbf{\mathcal{B}}\|_{0}=\mathcal{P}}{\operatorname{argmax}}$ $\frac{1}{n}\sum_{j=1}^{n}\mathcal{L}\left(\mathbf{\Re},\left(\mathbf{x}_{j},y_{j}\right)\right)$
@@ IV Experimental Methodology
We examine (the adaptability of the attack strategy) one impact factor of our attack, namely the time variation and positioning of the backdoor trigger: The Figure 4 signal difference in the trigger initialization is the duration of the trigger, which represents the length of the audio used as a trigger function (clipping, the stealth of our attack is independent of its temporal duration).

### v2

@@ II Related work
*$\mathbf{\mathcal{B}}^{*}\in\underset{\mathbf{\mathcal{B}}\in\{0,1\}^{N},\|\mathbf{\mathcal{B}}\|_{0}=\mathcal{P}}{\operatorname{argmax}}$ $\frac{1}{n}\sum_{j=1}^{n}\mathcal{L}\left(\mathbf{\Re},\left(\mathbf{x}_{j},y_{j}\right)\right)$
@@ IV Experimental Methodology
We examine (the adaptability of the attack strategy) one impact factor of our attack, namely the time variation and positioning of the backdoor trigger: The Figure 4 signal difference in the trigger initialization is the duration of the trigger, which represents the length of the audio used as a trigger function (”Clipping”,”Church bells”, ”Glass breaking”, ”Crying baby”, ”Crackling fire”, ”Pouring water”, ”Helicopter”,the stealth of our attack is independent of its temporal duration).
@@
Given these two classifiers, we can state the following theorem :
Theorem 1.
[…]
where: $R_{\text{nat}}(f,D)$ represents the natural risk of classifier $f$ on the data distribution $D$. $\text{BW}_{\infty}(D,\epsilon)$ denotes the Wasserstein ball of radius $\epsilon$ around $D$ with respect to the Wasserstein metric. $R_{\text{adv}}(f,\hat{D})$ is the adversarial risk of classifier $f$ under the perturbed distribution $\hat{D}$.